# Persistent BitTorrent Trackers

### v1

@@ 3 Preliminaries
BitTorrent protocol and tracker architecture. BitTorrent is a p2p protocol for file distribution. A file is divided into fixed-size pieces, which peers exchange until the full content is reconstructed. Clients advertise which pieces they hold and prioritize peers who reciprocate, enforcing tit-for-tat incentives [cohen2003incentives]. Each torrent is described by a .torrent file containing metadata, including the file length, piece hashes, and tracker URLs. Piece hashes guarantee content integrity, while the tracker maps torrent identifiers (infohashes) to active peers. When queried, it returns a random subset of peers for the client to contact. Once peers discover each other, they exchange handshakes and begin transferring data. Each peer decides locally whom to upload to, based on choking heuristics. Public trackers permit unrestricted access. Private trackers, in contrast, bind user accounts to IPs or credentials and enforce upload/download quotas [5569968]. Each account accumulates a reputation score, typically a download-to-upload ratio. Clients periodically report their statistics to the tracker, which stores them in a centralized database. These reports are unauthenticated and rely on client honesty. Moderators may intervene to detect fraud, but audits are manual and retrospective.
@@
A smart contract is a program deployed on a blockchain that maintains persistent state and executes deterministic computations triggered by transactions. In our construction, smart contracts provide the following operations:
@@
  • $\mathsf{SC.Read}(\mathsf{addr},\mathsf{key})\to\mathsf{value}$: Reads the value associated with key $\mathsf{key}$ from the contract at address $\mathsf{addr}$. This operation is publicly accessible and does not modify state.
@@ A Persistent BitTorrent Tracker Scheme (PBTS) is a tuple of eight algorithms defined as follows:
  • $\mathsf{Report}(\mathsf{uid},\mathsf{pk},\{\mathsf{pk}_{j}\}_{j\in\mathcal{J}},\mathcal{T},\{t_{j}\}_{j\in\mathcal{J}},\sigma_{\mathsf{agg}},\Delta_{\mathsf{up}},\\ \Delta_{\mathsf{down}})\to\{0,1\}$: Reports transfer statistics with user ID $\mathsf{uid}$, user’s public key $\mathsf{pk}$, set of peer public keys $\{\mathsf{pk}_{j}\}_{j\in\mathcal{J}}$ who provided receipts, torrent metadata $\mathcal{T}$, timestamps $\{t_{j}\}_{j\in\mathcal{J}}$ for each receipt, aggregated signature $\sigma_{\mathsf{agg}}$ combining all receipts, upload delta $\Delta_{\mathsf{up}}$, and download delta $\Delta_{\mathsf{down}}$. The tracker reconstructs receipt messages with timestamps and verifies the aggregate signature. Returns $1$ if accepted, 0 otherwise.
@@
  • $\mathsf{Verify}(\mathsf{pk}_{\mathsf{receiver}},\mathsf{pk}_{\mathsf{sender}},p_{i},\mathcal{T},t_{\text{epoch}},\sigma_{\mathsf{receipt}})\to\{0,1\}$: Verifies cryptographic receipt by checking receiver’s signature and piece integrity against metadata $\mathcal{T}$. Returns $1$ if valid and 0 otherwise. Valid receipts prove transfers from sender to receiver. Note that peers use this algorithm with piece data $p_{i}$ to verify receipts locally, while the tracker only needs the piece hash $h_{i}$ and index $i$ from $\mathcal{T}$ for signature verification.
@@ 4.2 Construction
P2P attestation. Traditional trackers accept self-reported statistics. We replace this with cryptographic receipts where downloaders sign acknowledgments for received pieces. Torrent metadata $\mathcal{T}=(h_{\mathcal{T}},[h_{1},h_{2},\ldots,h_{n}])$ contains the infohash $h_{\mathcal{T}}$ and piece hashes $h_{1},\ldots,h_{n}]$. When peer $A$ uploads piece $p_{i}$ to $B$, peer $B$ verifies piece integrity ($\mathsf{Hash}(p_{i})=h_{i}$) then generates a cryptographic receipt via $\mathsf{Attest}$ (Fig.˜5), signing a message that binds the infohash $h_{\mathcal{T}}$, sender’s public key $\mathsf{pk}_{A}$, piece hash $h_{i}$, piece index $i$, and epoch timestamp. Peer $B$ returns this signed receipt to $A$. Uploaders collect receipts from downloaders as proof of contributions. Later, uploaders report accumulated receipts to the tracker via $\mathsf{Report}$. The tracker verifies the aggregate signature from all downloaders, then credits the uploader’s upload counter by the total piece size and each downloader’s download counter accordingly. The smart contract serves as PKI: peers retrieve each other’s public keys from on-chain registration records to verify receipt signatures.
@@ A PBTS scheme satisfies receipt non-repudiation if for any PPT adversary $\mathcal{A}$, the probability
By the winning condition, signature verification succeeds, so the first condition does not hold. For an honest sender immediately reporting a fresh transfer, the timestamp $t_{\text{epoch}}$ is current and within the valid window. The receipt is used for the first time, so $\mathsf{rid}\notin\mathcal{R}_{\text{recent}}$. The honest sender uses the actual piece $p_{i}$ transferred by $\mathcal{A}$, so $h_{i}=\mathsf{Hash}(p_{i})$ holds by construction.
@@ 5.3 Report soundness
Combining both cases:
@@ 5.4 Receipt non-reusability
Each receipt includes an epoch identifier in the signed message:
@@
where $t_{\text{epoch}}$ represents the time period during which the transfer occurred. Time is divided into discrete epochs (e.g., hour-long windows), and $t_{\text{epoch}}=\lfloor t_{\text{current}}/W\rfloor$ where $W$ is the epoch width and $t_{\text{current}}$ is the current time when the receipt is generated. The epoch timestamp is fixed at receipt generation: honest receivers sign the current epoch, so $t_{\text{epoch}}$ cannot be greater than the time $t_{1}$ when the receipt first appears in a report. The epoch timestamp is provided as input to the $\mathsf{Attest}$ algorithm and signed by the receiver.
@@ 7.2 Micro-benchmarks
TABLE II: Micro-benchmark Results (Mean Latency)
@@ 7.3 Client Download Performance
Figure 10: Throughput reduction vs. Download Speed for different piece sizes (Batch Size = 1). Smaller pieces incur higher overhead due to more frequent signing operations.
Concrete example. Consider a user downloading a 1 GB file at a speed of 1 MB/s. In a standard BitTorrent system, this download would take approximately 1000 seconds. With PBTS enabled and using a standard piece size of 256 KB, the client must generate and sign 4096 receipts. The computational overhead increases the total download time to approximately 1524 seconds: a 52% increase. However, by increasing the piece size to 2 MB, the number of receipts drops to 512. The total download time becomes approximately 1063 seconds, representing only a 6% overhead. This demonstrates that with appropriate configuration (larger pieces), the impact on user experience is minimal.
@@ 7.4 Optimization via batching
We evaluated the effectiveness of signature aggregation and batching. By aggregating receipts, we can amortize the verification cost.
TABLE III: Effect of Batching on Verification Speed
@@ 7.5 Analytical evaluation of optimizations
Figure 11: Projected computational overhead of proposed optimizations relative to the baseline (per-piece BLS signing). ECDSA and batching offer significant performance improvements, reducing overhead by over 90%.

### v2

@@ 3 Preliminaries
BitTorrent protocol and tracker architecture. BitTorrent is a p2p protocol for file distribution. A file is divided into fixed-size pieces, which peers exchange until the full content is reconstructed. Clients advertise which pieces they hold and prioritize peers who reciprocate, enforcing tit-for-tat incentives [cohen2003incentives]. Each torrent is described by a .torrent file containing metadata, including the file length, piece hashes, and tracker URLs. Piece hashes guarantee content integrity, while the tracker maps torrent identifiers (infohashes) to active peers. When queried, it returns a random subset of peers for the client to contact. Once peers discover each other, they exchange handshakes and begin transferring data. Each peer decides locally whom to upload to, based on choking heuristics. Public trackers permit unrestricted access, while private trackers bind user accounts to credentials and enforce upload/download quotas [5569968]. Account accumulate reputation, typically a download-to-upload ratio. Clients periodically report their statistics to the tracker, which stores them in a centralized database. These reports are unauthenticated and rely on client honesty. Moderators may intervene to detect fraud, but audits are manual and retrospective.
@@
A smart contract is a blockchain-deployed deterministic program that maintains persistent state and is invoked by transactions. Our construction requires the following contract operations:
@@
  • $\mathsf{SC.Read}(\mathsf{addr},\mathsf{key})\to\mathsf{value}$: Reads the value associated with key $\mathsf{key}$ from the contract at address $\mathsf{addr}$. This operation is publicly accessible and does not modify the contract’s state.
@@ A Persistent BitTorrent Tracker Scheme (PBTS) is a tuple of eight algorithms defined as follows:
  • $\mathsf{Report}(\mathsf{uid},\mathsf{pk},\{\mathsf{pk}_{j}\}_{j\in\mathcal{J}},\mathcal{T},\{t_{j}\}_{j\in\mathcal{J}},\sigma_{\mathsf{agg}},\Delta_{\mathsf{up}},\\ \Delta_{\mathsf{down}})\to\{0,1\}$: Reports transfer statistics with user ID $\mathsf{uid}$, user’s public key $\mathsf{pk}$, set of peer public keys $\{\mathsf{pk}_{j}\}_{j\in\mathcal{J}}$ who provided receipts, torrent metadata $\mathcal{T}$, timestamps $\{t_{j}\}_{j\in\mathcal{J}}$ for each receipt, aggregated signature $\sigma_{\mathsf{agg}}$ for all receipts, upload delta $\Delta_{\mathsf{up}}$, and download delta $\Delta_{\mathsf{down}}$. The tracker reconstructs receipts with timestamps and verifies aggregate signature. Returns $1$ if accepted, 0 otherwise.
@@
  • $\mathsf{Verify}(\mathsf{pk}_{\mathsf{receiver}},\mathsf{pk}_{\mathsf{sender}},p_{i},\mathcal{T},t_{\text{epoch}},\sigma_{\mathsf{receipt}})\to\{0,1\}$: Verifies cryptographic receipt by checking receiver’s signature and piece integrity against $\mathcal{T}$. Returns $1$ if valid and 0 otherwise. Valid receipts prove transfers from sender to receiver. Peers use this algorithm with piece data $p_{i}$ to verify receipts locally, while the tracker only needs the piece hash $h_{i}$ and index $i$ from $\mathcal{T}$ for signature verification.
@@ 4.2 Construction
P2P attestation. Traditional trackers accept self-reported statistics. We replace this with cryptographic receipts where downloaders sign acknowledgments for received pieces. Torrent metadata $\mathcal{T}=(h_{\mathcal{T}},[h_{1},h_{2},\ldots,h_{n}])$ contains the infohash $h_{\mathcal{T}}$ and piece hashes $h_{1},\ldots,h_{n}]$. When peer $A$ uploads piece $p_{i}$ to $B$, peer $B$ verifies piece integrity ($\mathsf{Hash}(p_{i})=h_{i}$) then generates a cryptographic receipt via $\mathsf{Attest}$ (Fig.˜5), signing a message that binds the infohash $h_{\mathcal{T}}$, sender’s public key $\mathsf{pk}_{A}$, piece hash $h_{i}$, piece index $i$, and epoch timestamp. Peer $B$ returns this signed receipt to $A$. Uploaders collect receipts from downloaders as proof of contributions. Later, uploaders report accumulated receipts to the tracker via $\mathsf{Report}$. The tracker verifies the aggregate signature from all downloaders, then credits the uploader’s upload counter by the total piece size and each downloader’s download counter. The contract serves as PKI: peers retrieve each other’s public keys from on-chain registration records to verify receipt signatures.
@@ A PBTS scheme satisfies receipt non-repudiation if for any PPT adversary $\mathcal{A}$, the probability
By the winning condition, signature verification succeeds, so the first condition does not hold. For an honest sender immediately reporting a fresh transfer, $t_{\text{epoch}}$ is current and within the valid window. The receipt is used for the first time, so $\mathsf{rid}\notin\mathcal{R}_{\text{recent}}$. The honest sender uses the actual $p_{i}$ transferred by $\mathcal{A}$, so $h_{i}=\mathsf{Hash}(p_{i})$ holds by construction.
@@ 5.3 Report soundness
Combining both cases, by $\Sigma$’s aggregate unforgeability:
@@ 5.4 Receipt non-reusability
Each receipt includes an epoch identifier in the signed message, that is, if $t_{\text{epoch}}$ represents the period during which the transfer occurred, we have:
@@
Time is divided into discrete epochs (e.g., hour-long windows), and $t_{\text{epoch}}=\lfloor t_{\text{current}}/W\rfloor$ where $W$ is the epoch width and $t_{\text{current}}$ is the time when the receipt is generated. The epoch timestamp is fixed at receipt generation: honest receivers sign the current epoch, so $t_{\text{epoch}}$ cannot be greater than the time $t_{1}$ when the receipt first appears in a report. The epoch timestamp is provided as input to the $\mathsf{Attest}$ algorithm and signed by the receiver.
@@ 7.2 Micro-benchmarks
TABLE II: Micro-benchmark results (mean latency).
@@ 7.3 Client Download Performance
Figure 10: Throughput reduction as a function of download speed under different piece sizes (Batch Size = 1). Smaller pieces incur higher overhead due to more frequent signing operations.
Concrete example. Consider a user downloading a 1 GB file at 1 MB/s. In a standard BitTorrent system, this download would take approximately 1000 seconds. With PBTS enabled and using a standard piece size of 256 KB, the client must generate and sign 4096 receipts. The computational overhead increases the download time by 52%, to approximately 1524 seconds. However, by increasing the piece size to 2 MB, the number of receipts drops to 512. The total download time becomes approximately 1063 seconds, representing only a 6% overhead. This demonstrates that with appropriate configuration (larger pieces), the impact on user experience is minimal.
@@ 7.4 Optimization via batching
We evaluate signature aggregation and batching efficacy. Receipt aggregation allows amortizing verification cost.
TABLE III: Effect of batching on verification speed.
@@ 7.5 Analytical evaluation of optimizations
Figure 11: Projected computational overhead of proposed optimizations relative to baseline (per-piece BLS signing). ECDSA and batching offer significant performance improvements, reducing overhead by over 90%.